Formasa Plastics has major fabrication plants in Texas and Hong Kong. The president wants to know
the equivalent future worth of $1 million capital investments each year for 8 years, starting 1 year from
now. Formasa capital earns at a rate of 14% per year.
1
Expert's answer
2021-09-23T03:03:52-0400

Solution:

Future Worth of periodic payment is given by:

FW = (P/r) ((1 + r) n - 1)

Where P = Periodic amount = 1 million

 r = interest rate = 14% = 0.14

n = total number of periods = 8.

Hence, FW = (P/r) ((1 + r) n - 1) = (1 million/0.14) ((1 + 0.14)8 - 1)

Hence FW = $13232760.2 or (13.23 millions)

Hence Equivalent Future worth = $13232760.2 or (13.23 millions)